OpenStudy (anonymous):

Solve by factoring: 10t2 – t − 2 = 0

OpenStudy (saifoo.khan):

Look for two terms to whom, when you add you get -1 and when multiply -20

OpenStudy (anonymous):

10t^2 – t − 2 = 0 10t^2 -5t + 4t -2 =0 5t( 2t-1) + 2(t-1) =0 (2t-1)(5t+2) =0 (2t-1)=0 or (5t+2) =0 Can you do it from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am having difficulty with quadratics

OpenStudy (saifoo.khan):

Where are you stuck?

OpenStudy (anonymous):

I will go back and review. thanks for the help

OpenStudy (saifoo.khan):

I can guide you step by step if you stuck.. Sure.

OpenStudy (anonymous):

please it would be greatly appreciated.

OpenStudy (saifoo.khan):

Ok, so first step, look at the equation it says, : 10t2 – t − 2 = 0 a = 10 b = -1 c = -2 agree?

OpenStudy (anonymous):

yes correct

OpenStudy (saifoo.khan):

Ok, next step is to look for two terms.. First, when you add you get -1 and when multiply -20 [How i got -1? A: it's the value of b. How i got -20 it's the value of a * c. 10*-2 = -20 Agree?

OpenStudy (anonymous):

ok got that

OpenStudy (saifoo.khan):

So now two terms = -5 and 4 When we add them, we get -1 and when we multiply -20. agree?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Ok, so, 10t^2 – t − 2 = 0 break the middle term into two numbers we just figured.. 10t^2 -5t + 4t -2 =0

OpenStudy (anonymous):

ok got that

OpenStudy (saifoo.khan):

Now, follow @RolyPoly 's working. :)

OpenStudy (anonymous):

ok, thank you for helping me:)

OpenStudy (saifoo.khan):

No worries.